# Variation of cones of divisors in a family of varieties – Fano type case

### v3

@@ 5.1. Stable boundedness
${\rm{bc}}(X/T)\coloneqq\{h\mid h:X\dasharrow Y/T\text{~{}is a birational contraction}/T\text{~{}up to isomorphisms of~{}}Y/T\}.$
Note that the above set includes not only varieties birationally contracted by $X$ but also the birational contraction maps from $X$. However, for the sake of simplicity, we also write $Y/T\in{\rm{bc}}(X/T)$ instead of $[h:X\dasharrow Y/T]\in{\rm{bc}}(X/T)$. For $Y_{1}/T,Y_{2}/T\in{\rm{bc}}(X/T)$, $Y_{1}\simeq Y_{2}/T$ should be understood as that there exists an isomorphism $\theta:Y_{1}\to Y_{2}$, such that $h_{2}=\theta\circ h_{1}:X\dasharrow Y_{2}$, where $h_{i}:X\dasharrow Y_{i},i=1,2$ are given birational contractions. Moreover, for a subvariety $V\subset T$, we write $Y_{V}/V$ instead of the base change of $h$ to $V$.
@@
a partial MMP$/T$ (with respect to a divisor $D$) if each step consists of $D_{i}$-extremal divisorial/flipping contractions and $D_{i}$-flips, where $D_{i}$ is the strict transform of $D$ on $X_{i}$. Note that $D_{n}$ is not required to be nef over $T$. On the other hand, if $D_{n}$ becomes nef and big over $T$, and $X_{n}\to X_{n+1}$ is the birational contraction induced by $D_{n}$, then
@@
is also called a partial MMP$/T$. We set
$\begin{split}{\rm{bc}}(X/T)_{\rm MMP}\coloneqq\{h\mid&X\dasharrow Y/T\text{~{}is a partial~{}}\text{MMP}/T\\ &\text{up to isomorphisms of~{}}Y/T\}.\end{split}$ (5.1)
@@
Step 3. Let ${\rm{Exc}}(h_{1})\subset Y_{V}^{(1)}$ be the exceptional divisor. We also denote the corresponding divisor on $Y^{(1)}$ by ${\rm{Exc}}(h_{1})$. Note that ${\rm{Exc}}(h_{1})\cap Y^{(1)}_{\Omega}=\emptyset$. Let $Y^{(2)}\to Y^{(1)}$ be a small $\mathbb{Q}$-factorial modification. Let $E^{(2)}$ be the strict transform of ${\rm{Exc}}(h_{1})$. Let $Y^{(2)}\dasharrow Y^{(3)}$ be the canonical model of $E^{(2)}$ over $Y^{(1)}$. In particular, if $E^{(3)}$ is the strict transform of $E^{(2)}$, then $E^{(3)}$ is a $\mathbb{Q}$-Cartier divisor. Note that we still have $Y^{(3)}_{\Omega}\simeq Y^{(1)}_{\Omega}\simeq Y^{\prime}_{\Omega}$, and $Y^{(3)}_{V}\dasharrow Y^{\prime}$ is a birational contraction.
Step 4. We want to modify $Y^{(3)}$ to $Y^{(4)}$ so that $h_{4}:Y^{(4)}_{V}\dasharrow Y^{\prime}$ is an isomorphism in codimension $1$. Let ${\rm{Exc}}(h_{3})\subset Y_{V}^{(3)}$ be the exceptional divisor, where $h_{3}:Y_{V}^{(3)}\dasharrow Y^{\prime}$. We denote the corresponding divisor on $Y^{(3)}$ by ${\rm{Exc}}_{3}$. By construction, ${\rm{Exc}}(h_{3})$ is vertical over $V$. We claim that ${\rm{Exc}}(h_{3})$ is a very exceptional divisor over $V$. Suppose otherwise, by the smoothness of $V$, there exists a divisor $D>0$ such that ${\rm{Supp}}(f_{3}^{*}D)\subset{\rm{Exc}}(h_{3})$, where $f_{3}:Y^{(3)}_{V}\to V$. In particular, $f^{\prime*}D$ is a divisor on $Y^{\prime}$, where $f^{\prime}:Y^{{}^{\prime}}_{V}\to V$. However, any irreducible component of ${\rm{Supp}}(f^{\prime*}D)$ cannot belong to ${\rm{Supp}}(f_{3}^{*}D)$ as ${\rm{Supp}}(f_{3}^{*}D)\subset{\rm{Exc}}(h_{3})$ is an exceptional divisor. This means that $h_{3}^{-1}$ must contract ${\rm{Supp}}(f^{\prime*}D)$, and thus contradicts that $h_{3}$ is a birational contraction. As the center of each component of ${\rm{Exc}}_{3}$ lies in $V$, we see that ${\rm{Exc}}_{3}$ is also a very exceptional divisor on $Y^{(3)}$ over $T_{0}$. Runing an ${\rm{Exc}}_{3}$-MMP/$T_{0}$, $Y^{(3)}\dasharrow Y^{(4)}$, we can contract ${\rm{Exc}}_{3}$. Moreover, we still have $Y^{(4)}_{\Omega}\simeq Y^{(3)}_{\Omega}\simeq Y^{\prime}_{\Omega}$. This shows that $h_{4}:Y^{(4)}_{V}\dasharrow Y^{\prime}$ is an isomorphism in codimension $1$.
Step 5. Let $H^{\prime}$ be a general very ample$/V$ divisor on $Y^{\prime}$. Let $H^{(4)}$ be the strict transform of $H^{\prime}$ on $Y^{(4)}$. Note that $H^{(4)}$ may not be $\mathbb{Q}$-Cartier. Let $Y^{(5)}\to Y^{(4)}$ be a small $\mathbb{Q}$-factorial modification, and let $H^{(5)}$ be the strict transform of $H^{\prime}$ on $Y^{(5)}$. Let $Y^{(5)}\dasharrow Y^{(6)}$ be the canonical model of $H^{(5)}$ over $T_{0}$. Then, $H^{(6)}$, as the strict transform of $H^{\prime}$ on $Y^{(5)}$, is ample over $T_{0}$. Moreover, we have $Y^{(6)}_{\Omega}\simeq Y^{\prime}_{\Omega}$ by construction. As $h_{5}:Y^{(5)}_{V}\dasharrow Y^{\prime}$ is an isomorphism in codimension $1$, and $h_{6}:Y^{(6)}_{V}\dasharrow Y^{\prime}$ is an isomorphism on the locus where $h_{5}$ is an isomorphism, we see that $h_{6}$ is still an isomorphism in codimension $1$.
@@ 5.2. MMP in a family
${\rm{bc}}(X/T)_{\rm MMP}\coloneqq\{h\mid X\dasharrow Y/T\text{~{}is a partial~{}}\text{MMP}/T\text{~{}up to isomorphisms of~{}}Y/T\}$
@@
    Let $X\dashrightarrow Y/T$ be a partial MMP$/T$ consisting of divisorial contractions, flipping contractions and flips. Suppose that the natural map $N^{1}(X/T)\to N^{1}(X_{t})$ is surjective for very general $t\in T$. Then there exists an open subset $T_{0}\subset T$ such that the natural map
@@
Step 1. First, we prove statement (1). Shrinking $T$, we can assume that each step of the above MMP$/T$ is not a constant map over the generic point of $T$. We claim that the desired property is preserved in each step of this MMP, and thus it holds for $Y/T$.
@@
Suppose that $h:Z\to Y^{\prime}/T$ is a birational contraction between varieties of Fano type$/T$ and $N^{1}(Z/T)\to N^{1}(Z_{t})$ is surjective for very general $t\in T$. Note that $h$ corresponds to the contraction of curves belonging to a face $F$ of the Mori cone ${\rm{NE}}(Z/Y^{\prime})$. We can contract an extremal ray of $F$ to obtain a morphism $h^{\prime}:Z\to Z^{\prime}$ and thus $h$ factors through $h^{\prime}$. By repeating this process, we can further assume that $h$ is an extremal contraction. Then the desired property follows from Step 1. ∎
@@
Let $h:Y\to X$ be a log resolution of $(X,\Delta)$. Shrinking $T$, we can assume that $Y\to T$ is a smooth morphism. As $(X,\Delta)$ has klt singularities, after shrinking $T$, we can assume that $(X_{s},\Delta_{s})$ still has klt singularties for $s\in S$. In particular, $X_{s}$ has rational singularities (see [KM98, Theorem 5.22]). Thus, possibly shrinking $T$ further, $H^{i}(Y_{s},\mathcal{O}_{Y_{s}})=0,i=1,2$ for each $s\in S$. By Theorem 1.3 (2), there exits a generic finite base change $T^{\prime}\to T$ and an open subset $T^{\prime}_{0}\subset T^{\prime}$ such that $N^{1}(X_{T_{0}^{\prime}}/T_{0}^{\prime})\to N^{1}(X_{t^{\prime}})$ is surjective for each $t^{\prime}\in T_{0}^{\prime}$. Shrinking $T_{0}^{\prime}$, we can assume that $T_{0}^{\prime}$ is a smooth variety. Thus $Y_{T^{\prime}_{0}}$ is still a smooth variety. Applying Theorem 1.3 (2) to $Y_{T^{\prime}_{0}}\to T^{\prime}_{0}$, there exits a generic finite base change $T^{\prime\prime}\to{T^{\prime}_{0}}$ and an open subset $T_{0}^{\prime\prime}\subset T^{\prime\prime}$ such that $N^{1}(Y_{T_{0}^{\prime\prime}}/T_{0}^{\prime\prime})\to N^{1}(Y_{t^{\prime\prime}})$ is surjective for each $t^{\prime\prime}\in T_{0}^{\prime\prime}$. Note that we still have the surjectivity of the map $N^{1}(X_{T_{0}^{\prime\prime}}/T_{0}^{\prime\prime})\to N^{1}(X_{t^{\prime\prime}}),t^{\prime\prime}\in T_{0}^{\prime\prime}$. Shrinking $T_{0}^{\prime\prime}$, we can assume that there still exists a $\mathbb{Q}$-divisor $\Delta^{\prime\prime}$ such that $(X_{T_{0}^{\prime\prime}},\Delta^{\prime\prime})$ has klt singularities. Thus, we can replace $T^{\prime\prime}_{0}$ by $T$ and assume that
@@
Now we proceed with the standard process of small $\mathbb{Q}$-factorial modifications. Let $K_{Y}+\Delta_{Y}=h^{*}(K_{X}+\Delta)+E$ such that ${\rm{Supp}}E={\rm{Supp}}{\rm{Exc}}(h)$ and $(Y,\Delta_{Y})$ has klt singularities. Shrinking $T$, we can assume that $h$ is a fiber-wise resolution such that the same property of $K_{Y_{t}}+\Delta_{Y_{t}}=h_{t}^{*}(K_{X_{t}}+\Delta_{t})+E_{t}$ holds fiber-wise, where $\Delta_{Y_{t}}=(\Delta_{Y})|_{Y_{t}}$. By [BCHM10], there exists a $(K_{Y}+\Delta_{Y})$-MMP over $X$ which terminates and only $E$ is contracted in this process. Denote this MMP$/X$ by
@@
For the inverse inclusion, let $Y/T\in{\rm{bc}}(X/T)$. We show that $Y$ can be obtained from $X$ via an MMP$/T$ followed by a birational contraction. To be precise, we need to show that the birational contraction $X\dashrightarrow Y/T$ corresponding to $Y$ can be factored into an MMP$/T$ followed by a birational contraction up to an isomorphism of $Y/T$.
Let $A^{\prime}$ be an ample$/T$ divisor on $Y$, and let $A$ be its strict transform on $X$. Let $p:W\to X$ and $q:W\to Y$ be birational morphisms such that $q\circ p^{-1}$ is the map $X\dashrightarrow Y$. As $X$ is a $\mathbb{Q}$-factorial variety, we can write
@@
Since $A^{\prime}$ is ample$/T$ and both $F$ and $E_{2}$ are $q$-exceptional divisors, the divisors $q^{*}A^{\prime}$ and $F+E_{2}$ form the positive and negative parts, respectively, of the Nakayama-Zariski decomposition of $p^{*}B+E_{1}$ over $T$. Let $X\dashrightarrow Z/T$ be a $B$-MMP$/T$ such that $B_{Z}$, the strict transform of $B$, is semi-ample$/T$. Replacing $W$ by a higher model, we can assume that there exists a birational morphism $r:W\to Z$. We have
@@
where $E_{Z}$ is an effective $r$-exceptional divisor. Since $E_{1}$ is also $r$-exceptional, $p^{*}B+E_{1}$ admits the Nakayama-Zariski decomposition whose positive part is $r^{*}B_{Z}$ and the negative part is $E_{Z}+E_{1}$. By the uniqueness of the Nakayama-Zariski decomposition, we have $E_{Z}+E_{1}=F+E_{2}$ and $r^{*}B_{Z}=q^{*}A^{\prime}$. As $A^{\prime}$ is ample$/T$, the natural birational map $g:Z\dashrightarrow Y/T$ is indeed a morphism. ∎
@@
Let $X$ be of Fano type over $T$. Then, up to a generic finite base change of $T$, there exists an open subset $T_{0}\subset T$ such that the following properties hold:
@@
    for any $Y/T_{0}\in{\rm{bc}}(X/T_{0})$, $Y$ is flat over $T_{0}$, and $Y_{t}$ is an irreducible normal variety for each $t\in T_{0}$.
@@
By Theorem 1.4, there exists a generic finite base change $T^{\prime}\to T$ and an open subset $T_{0}^{\prime}\subset T^{\prime}$ such that $X_{T_{0}^{\prime}}$ admits a small $\mathbb{Q}$-factorial modification $W\to X_{T_{0}^{\prime}}$ which is also a fiber-wise small $\mathbb{Q}$-factorial modification over $T_{0}^{\prime}$. Moreover, the natural map $N^{1}(W/U)\to N^{1}(W_{t}),t\in U$ is an isomorphism, where $U\subset T_{0}^{\prime}$ is an open subset. Note that $W$ is of Fano type over $T_{0}^{\prime}$ after shrinking $T_{0}^{\prime}$. By Proposition 5.4, we have
@@
For (2), after shrinking $T$, we can assume that for any $Y/T\in{\rm{bc}}(X/T)$, $Y$ is flat over $T$ by generic flatness, and $Y_{t},t\in T$ is an irreducible normal variety as $\{p\in T\mid Y_{p}\text{~{}is normal}\}$ is a constructible set. Note that, in the above set, $p$ is not restricted to closed points, and the generic point of $T$ belongs to this set as $Y$ is normal.
@@
By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. After replacing $T$ by a generic finite base change and possibly shrinking it, we may assume that the properties listed in Lemma 5.5 hold.
First, we show that an MMP of the total space is an MMP of each fiber of the same type. For any $\mathcal{D}\in N^{1}(X/T)$, let
@@
be a $\mathcal{D}$-MMP over $T$, where $X_{i}\dasharrow X_{i+1},i=0,\ldots,n-1$ are birational contractions, and $X_{n}\to X_{n+1}$ is either a contraction induced by the semi-ample$/T$ divisor $\mathcal{D}_{X_{n}}$ or a Mori fiber space when $\mathcal{D}$ is not pseudo-effective$/T$. Thus, we have $X_{i}/T\in{\rm{bc}}(X/T)_{\rm MMP},0\leq i\leq n$, and $X_{n+1}/T\in{\rm{bc}}(X/T)_{\rm MMP}$ if $X\dasharrow X_{n+1}$ is birational. To be precise, this means that the above natural birational contraction $X\dasharrow X_{i}/T$ belongs to ${\rm{bc}}(X/T)_{\rm MMP}$. By Lemma 5.5 (3), if $g:X_{i}\dasharrow X_{i+1}$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space) if and only if $g_{t}$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space) for each $t\in T$. Hence, for each $t\in T$,
@@
Conversely, we show that an MMP of the fiber is induced from an MMP$/T$ of the total space. Suppose that

### v4

@@ 5.1. Stable boundedness
[…]
Note that the above set includes not only varieties birationally contracted by $X$ but also the birational contraction maps from $X$. However, for the sake of simplicity, we also write $Y/T\in{\rm{bc}}(X/T)$ instead of $h:X\dasharrow Y/T]\in{\rm{bc}}(X/T)$. For $Y_{1}/T,Y_{2}/T\in{\rm{bc}}(X/T)$, $Y_{1}\simeq Y_{2}/T$ should be understood as that there exists an isomorphism $\theta:Y_{1}\to Y_{2}$, such that $h_{2}=\theta\circ h_{1}:X\dasharrow Y_{2}$, where $h_{i}:X\dasharrow Y_{i},i=1,2$ are given birational contractions. Moreover, for a subvariety $V\subset T$, we write $Y_{V}/V$ instead of the base change of $h$ to $V$.
@@
a partial MMP/T/T (with respect to a divisor $D$) if each step consists of $D_{i}$-extremal divisorial/flipping contractions and $D_{i}$-flips, where $D_{i}$ is the strict transform of $D$ on $X_{i}$. Note that $D_{n}$ is not required to be nef over $T$. On the other hand, if $D_{n}$ becomes nef and big over $T$, and $X_{n}\to X_{n+1}$ is the birational contraction induced by $D_{n}$, then
@@
is also called a partial MMP/T/T. We set
[…]
@@
Step 3. Let ${\rm{Exc}}(h_{1})\subset Y_{V}^{(1)}$ be the exceptional divisor. We also denote the corresponding divisor on $Y^{(1)}$ by ${\rm{Exc}}(h_{1})$. Note that ${\rm{Exc}}(h_{1})\cap Y^{(1)}_{\Omega}$$\emptyset$. Let $Y^{(2)}\to Y^{(1)}$ be a small $\mathbb{Q}$-factorial modification. Let $E^{(2)}$ be the strict transform of ${\rm{Exc}}(h_{1})$. Let $Y^{(2)}\dasharrow Y^{(3)}$ be the canonical model of $E^{(2)}$ over $Y^{(1)}$. In particular, if $E^{(3)}$ is the strict transform of $E^{(2)}$, then $E^{(3)}$ is a $\mathbb{Q}$-Cartier divisor. Note that we still have $Y^{(3)}_{\Omega}\simeq Y^{(1)}_{\Omega}\simeq Y^{\prime}_{\Omega}$, and $Y^{(3)}_{V}\dasharrow Y^{\prime}$ is a birational contraction.
[…]
Step 5. Let $H^{\prime}$ be a general very ample/V/V divisor on $Y^{\prime}$. Let $H^{(4)}$ be the strict transform of $H^{\prime}$ on $Y^{(4)}$. Note that $H^{(4)}$ may not be $\mathbb{Q}$-Cartier. Let $Y^{(5)}\to Y^{(4)}$ be a small $\mathbb{Q}$-factorial modification, and let $H^{(5)}$ be the strict transform of $H^{\prime}$ on $Y^{(5)}$. Let $Y^{(5)}\dasharrow Y^{(6)}$ be the canonical model of $H^{(5)}$ over $T_{0}$. Then, $H^{(6)}$, as the strict transform of $H^{\prime}$ on $Y^{(5)}$, is ample over $T_{0}$. Moreover, we have $Y^{(6)}_{\Omega}\simeq Y^{\prime}_{\Omega}$ by construction. As $h_{5}:Y^{(5)}_{V}\dasharrow Y^{\prime}$ is an isomorphism in codimension $1$, and $h_{6}:Y^{(6)}_{V}\dasharrow Y^{\prime}$ is an isomorphism on the locus where $h_{5}$ is an isomorphism, we see that $h_{6}$ is still an isomorphism in codimension $1$.
@@ 5.2. MMP in a family
[…]
@@
    Let $X\dashrightarrow Y/T$ be a partial MMP/T/T consisting of divisorial contractions, flipping contractions and flips. Suppose that the natural map $N^{1}(X/T)\to N^{1}(X_{t})$ is surjective for very general $t\in T$. Then there exists an open subset $T_{0}\subset T$ such that the natural map
@@
Step 1. First, we prove statement (1). Shrinking $T$, we can assume that each step of the above MMP/T/T is not a constant map over the generic point of $T$. We claim that the desired property is preserved in each step of this MMP, and thus it holds for $Y/T$.
@@
Suppose that $h:Z\to Y^{\prime}/T$ is a birational contraction between varieties of Fano type/T/T and $N^{1}(Z/T)\to N^{1}(Z_{t})$ is surjective for very general $t\in T$. Note that $h$ corresponds to the contraction of curves belonging to a face $F$ of the Mori cone ${\rm{NE}}(Z/Y^{\prime})$. We can contract an extremal ray of $F$ to obtain a morphism $h^{\prime}:Z\to Z^{\prime}$ and thus $h$ factors through $h^{\prime}$. By repeating this process, we can further assume that $h$ is an extremal contraction. Then the desired property follows from Step 1. ∎
@@
Let $h:Y\to X$ be a log resolution of $(X,\Delta)$. Shrinking $T$, we can assume that $Y\to T$ is a smooth morphism. As $X_{s},s\in S$ has rational singularities with $h^{i}(X_{s},\mathcal{O}_{X_{s}})=0,i=1,2$, we can assume that the same property still holds for $Y_{s},s\in S$ after shrinking $T$. By Theorem 1.3 (2), there exits a generically finite base change $T^{\prime}\to T$ and an open subset $T^{\prime}_{0}\subset T^{\prime}$ such that $N^{1}(X_{T_{0}^{\prime}}/T_{0}^{\prime})\to N^{1}(X_{t^{\prime}})$ is surjective for each $t^{\prime}\in T_{0}^{\prime}$. Shrinking $T_{0}^{\prime}$, we can assume that $T_{0}^{\prime}$ is a smooth variety. Thus $Y_{T^{\prime}_{0}}$ is still a smooth variety. Applying Theorem 1.3 (2) to $Y_{T^{\prime}_{0}}\to T^{\prime}_{0}$, there exits a generically finite base change $T^{\prime\prime}\to{T^{\prime}_{0}}$ and an open subset $T_{0}^{\prime\prime}\subset T^{\prime\prime}$ such that $N^{1}(Y_{T_{0}^{\prime\prime}}/T_{0}^{\prime\prime})\to N^{1}(Y_{t^{\prime\prime}})$ is surjective for each $t^{\prime\prime}\in T_{0}^{\prime\prime}$. Note that we still have the surjectivity of the map $N^{1}(X_{T_{0}^{\prime\prime}}/T_{0}^{\prime\prime})\to N^{1}(X_{t^{\prime\prime}}),t^{\prime\prime}\in T_{0}^{\prime\prime}$. Shrinking $T_{0}^{\prime\prime}$, we can assume that there still exists a $\mathbb{Q}$-divisor $\Delta^{\prime\prime}$ such that $(X_{T_{0}^{\prime\prime}},\Delta^{\prime\prime})$ has klt singularities. Thus, we can replace $T^{\prime\prime}_{0}$ by $T$ and assume that
@@
Now we proceed with the standard process of small $\mathbb{Q}$-factorial modifications. Let $K_{Y}+\Delta_{Y}=h^{*}(K_{X}+\Delta)+E$ such that ${\rm{Supp}}E={\rm{Supp}}{\rm{Exc}}(h)$ and $(Y,\Delta_{Y})$ has klt singularities. Shrinking $T$, we can assume that $h$ is a fiber-wise resolution such that the same property of $K_{Y_{t}}+\Delta_{Y_{t}}=h_{t}^{*}(K_{X_{t}}+\Delta_{t})+E_{t}$ holds fiber-wise, where $\Delta_{Y_{t}}=(\Delta_{Y})|_{Y_{t}}$. By [BCHM10], there exists a $(K_{Y}+\Delta_{Y})$-MMP over $X$ which terminates and only $E$ is contracted in this process. Denote this MMP/X/X by
@@
For the inverse inclusion, let $Y/T\in{\rm{bc}}(X/T)$. We show that $Y$ can be obtained from $X$ via an MMP/T/T followed by a birational contraction. To be precise, we need to show that the birational contraction $X\dashrightarrow Y/T$ corresponding to $Y$ can be factored into an MMP/T/T followed by a birational contraction up to an isomorphism of $Y/T$.
Let $A^{\prime}$ be an ample/T/T divisor on $Y$, and let $A$ be its strict transform on $X$. Let $p:W\to X$ and $q:W\to Y$ be birational morphisms such that $q\circ p^{-1}$ is the map $X\dashrightarrow Y$. As $X$ is a $\mathbb{Q}$-factorial variety, we can write
@@
Since $A^{\prime}$ is ample/T/T and both $F$ and $E_{2}$ are $q$-exceptional divisors, the divisors $q^{*}A^{\prime}$ and $F+E_{2}$ form the positive and negative parts, respectively, of the Nakayama-Zariski decomposition of $p^{*}B+E_{1}$ over $T$. Let $X\dashrightarrow Z/T$ be a $B$-MMP/T/T such that $B_{Z}$, the strict transform of $B$, is semi-ample/T/T. Replacing $W$ by a higher model, we can assume that there exists a birational morphism $r:W\to Z$. We have
@@
where $E_{Z}$ is an effective $r$-exceptional divisor. Since $E_{1}$ is also $r$-exceptional, $p^{*}B+E_{1}$ admits the Nakayama-Zariski decomposition whose positive part is $r^{*}B_{Z}$ and the negative part is $E_{Z}+E_{1}$. By the uniqueness of the Nakayama-Zariski decomposition, we have $E_{Z}+E_{1}=F+E_{2}$ and $r^{*}B_{Z}=q^{*}A^{\prime}$. As $A^{\prime}$ is ample/T/T, the natural birational map $g:Z\dashrightarrow Y/T$ is indeed a morphism. ∎
@@
Let $X$ be of Fano type over $T$. Then, up to a generically finite base change of $T$, there exists an open subset $T_{0}\subset T$ such that the following properties hold:
@@
    for any $T_{0}\in{\rm{bc}}(X/T_{0}), YY$ is flat over $T_{0}$, and $Y_{t}$ is an irreducible normal variety for each $t\in T_{0}$.
@@
By Theorem 1.4, there exists a generically finite base change $T^{\prime}\to T$ and an open subset $T_{0}^{\prime}\subset T^{\prime}$ such that $X_{T_{0}^{\prime}}$ admits a small $\mathbb{Q}$-factorial modification $W\to X_{T_{0}^{\prime}}$ which is also a fiber-wise small $\mathbb{Q}$-factorial modification over $T_{0}^{\prime}$. Moreover, the natural map $N^{1}(W/U)\to N^{1}(W_{t}),t\in U$ is an isomorphism, where $U\subset T_{0}^{\prime}$ is an open subset. Note that $W$ is of Fano type over $T_{0}^{\prime}$ after shrinking $T_{0}^{\prime}$. By Proposition 5.4, we have
@@
For (2), after shrinking $T$, we can assume that for any $T\in{\rm{bc}}(X/T), YY$ is flat over $T$ by generic flatness, and $Y_{t},t\in T$ is an irreducible normal variety as $\{p\in T\mid Y_{p}\text{~is normal}\}$ is a constructible set. Note that, in the above set, $p$ is not restricted to closed points, and the generic point of $T$ belongs to this set as $Y$ is normal.
@@
By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. After replacing $T$ by a generically finite base change and possibly shrinking it, we may assume that the properties listed in Lemma 5.5 hold.
First, by Proposition 5.1 and Proposition 5.2, it suffices to consider a fixed birational contraction $g:X^{\prime}\dashrightarrow Y/U$. Take birational morphisms $p\colon W\to X^{\prime}$ and $q\colon W\to Y$ such that $g=q\circ p^{-1}$. Shrinking $U$, we may assume that for any $t\in U$, the induced map
$g_{t}\colon X^{\prime}_{t}\dashrightarrow Y_{t}$
is still a birational contraction, and that $p_{t}$ and $q_{t}$ are birational morphisms. Moreover, we may further assume that if a divisor $E$ on $W$ is $p$-exceptional (resp. $q$-exceptional), then for every $t\in U$, the divisor $E|_{W_{t}}$ is $p_{t}$-exceptional (resp. $q_{t}$-exceptional). By construction, for any $\mathbb{R}$-Cartier divisor $\mathcal{D}$ on $X^{\prime}$, we have
$p^{*}\mathcal{D}=q^{*}(g_{*}\mathcal{D})+E,$
where $E$ is $q$-exceptional. Hence, we have
$p_{t}^{*}\mathcal{D}|_{X_{t}^{\prime}}\sim_{\mathbb{R}}q_{t}^{*}((g_{*}\mathcal{D})|_{Y_{t}})+E|_{W_{t}}.$
As $g_{t}$ is a birational contraction and $E|_{W_{t}}$ is $q_{t}$-exceptional, we see
$g_{t,*}(\mathcal{D}|_{X_{t}^{\prime}})\sim_{\mathbb{R}}(g_{*}\mathcal{D})|_{Y_{t}}.$
Next, we show that an MMP of the total space is an MMP of each fiber of the same type. For any $\mathcal{D}\in N^{1}(X/T)$, let
@@
be a $\mathcal{D}$-MMP over $T$, where $X_{i}\dasharrow X_{i+1},i=0,\ldots,n-1$ are birational contractions, and $X_{n}\to X_{n+1}$ is either a contraction induced by the semi-ample/T/T divisor $\mathcal{D}_{X_{n}}$ or a Mori fiber space when $\mathcal{D}$ is not pseudo-effective/T/T. Thus, we have $X_{i}/T\in{\rm{bc}}(X/T)_{\rm MMP},0\leq i\leq n$, and $X_{n+1}/T\in{\rm{bc}}(X/T)_{\rm MMP}$ if $X\dasharrow X_{n+1}$ is birational. To be precise, this means that the above natural birational contraction $X\dasharrow X_{i}/T$ belongs to ${\rm{bc}}(X/T)_{\rm MMP}$. By Lemma 5.5 (3), if $g:X_{i}\dasharrow X_{i+1}$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space) if and only if $g_{t}$ is a divisorial contraction (resp. a small contraction, an extremal contraction, a Mori fiber space) for each $t\in T$. Hence, for each $t\in T$,
@@
Conversely, we show that an MMP of the fiber is induced from an MMP/T/T of the total space. Suppose that